# Computing a Fixed Point of Contraction Maps in Polynomial Queries

### v1

@@ Abstract
We give an algorithm for finding an $\varepsilon$-fixed point of a contraction map $f:[0,1]^{k}\mapsto[0,1]^{k}$ under the $\ell_{\infty}$-norm with query complexity $O(k^{2}\log(1/\varepsilon))$.
@@ 1 Introduction
A particularly important metric space to study the problem of finding a Banach’s fixed point is the $k$-cube $[0,1]^{k}$ with respect to the $\ell_{\infty}$-norm, since many important problems can be reduced to that of finding an $\varepsilon$-fixed point (i.e., $x\in[0,1]^{k}$ satisfying $\|f(x)-x\|_{\infty}\leq\varepsilon$) in a ${(1-\gamma)}$-contraction map under the $\ell_{\infty}$-norm. Such problems arise from a variety of fields including stochastic analysis, optimization, verification, semantics, and game theory. For example, the classical dynamic programming models mentioned above (Markov decision processes etc.) involve contraction maps under the $\ell_{\infty}$-norm. Furthermore, the same holds for several well-known open problems that have been studied extensively and are currently not known to be in ${\mathsf{P}}$. For instance, Condon’s simple stochastic games (SSGs) [Con92] can be reduced to the problem of finding an $\varepsilon$-fixed point in a $(1-\gamma)$-contraction map over $[0,1]^{k}$ under the $\ell_{\infty}$-norm. A similar reduction from [EY10] extends to an even broader class of games, namely, Shapley’s stochastic games [Sha53], which lay the foundation of multi-agent reinforcement learning [Lit94]. The same holds also of course for other problems known to be subsumed by SSGs, like parity games, which are important in verification (see e.g. [EJ91, CJK+22]), and mean payoff games [ZP96]. Crucially, in all these reductions, both the approximation parameter $\varepsilon$ and the contraction parameter $\gamma$ are inversely exponential in the input size. Therefore, efficient algorithms in this context are those with a complexity upper bound that is polynomial in $k,\log(1/\varepsilon)$ and $\log(1/\gamma)$.
In this paper we consider general algorithms that access the contraction map in a black-box manner (as an oracle), and study the query complexity of finding an $\varepsilon$-fixed point of a $(1-\gamma)$-contraction map over the $k$-cube $[0,1]^{k}$ under the $\ell_{\infty}$-norm (which we denote by $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$). An algorithm under this model is given $k$, $\varepsilon$, $\gamma$, and oracle access to an unknown $(1-\gamma)$-contraction map $f$ over $[0,1]^{k}$. In each round the algorithm can send a point $x\in[0,1]^{k}$ to the oracle to reveal its value $f(x)$. The goal of the algorithm is to find an $\varepsilon$-fixed point with as few queries as possible.
@@
There is an $O(k^{2}\log({1}/{\varepsilon}))$-query algorithm for $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$.
@@
$\varepsilon/2\geq\|g(x)-x\|_{\infty}=\|(1-\varepsilon/2)f(x)-x\|_{\infty}\geq\|f(x)-x\|_{\infty}-\varepsilon/2$
@@
In Section 4, we give an $O(k^{2}\log(1/\varepsilon))$-query algorithm for $\textsc{Contraction}_{\infty}(\varepsilon/2,\varepsilon/2,k)$, from which Theorem 1 follows. Indeed, note that Observation 1 holds even if $f$ is a non-expansive map (i.e., $f$ has Lipschitz constant $1$: $\|f(x)-f(y)\|_{\infty}\leq\|x-y\|_{\infty}$ for all $x,y\in[0,1]^{k}$). As a result, the same query upper bound applies to $\textsc{NonExp}_{\infty}(\varepsilon,k)$, the problem of finding an $\varepsilon$-fixed point in a non-expansive map over $[0,1]^{k}$ under the $\ell_{\infty}$-norm:
@@
There is an $O(k^{2}\log({1}/{\varepsilon}))$-query algorithm for $\textsc{NonExp}_{\infty}(\varepsilon,k)$.
@@
There is an $O(k^{2}\log({1}/({\varepsilon\gamma})))$-query algorithm for $\textsc{StrContraction}_{\infty}(\varepsilon,\gamma,k)$.
In sharp contrast with Corollary 2, however, we show that it is impossible to strongly approximate an exact fixed point in a non-expansive map over $[0,1]^{2}$ under the $\ell_{\infty}$ norm.
@@
[…]
@@
[…]
@@
Remark. It is also important to note that while our algorithm in Theorem 1 is query efficient, it is not time efficient for the current version. The algorithm guarantees that within polynomial queries we can find a weak $\varepsilon$-fixed point, but each iteration requires a brute force procedure to determine the next query point. We will explain more details of techniques in Section 1.1.
@@
Other Related Work. We have already mentioned the most relevant works addressing the query complexity of computing the fixed point of a contraction map. For continuous functions $f:[0,1]^{k}\mapsto[0,1]^{k}$ that have Lipschitz constant greater than 1 (i.e. are expansive), there are exponential lower bounds on the query complexity of computing a (weak) approximate fixed point [HPV89, CD08].
@@
$\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$ when considered in the white-box model [^4] The white-box model refers to the model where the function is explicitly given by a polynomial-size circuit, in contrast to the black-box model where the function can only be accessed via an oracle as we studied in this paper. When we talk about a computational problem under the white-box model, we measure the efficiency by the time complexity. can be formulated as a total search problem so that it lies in the class ${\mathsf{TFNP}}$. In fact, it is one of the motivating problems in [DP11] to define the class CLS for capturing problems that lie in both ${\mathsf{PLS}}$ [JPY88] and ${\mathsf{PPAD}}$ [Pap94]. Later, it is placed in UEOPL [FGMS20] [^5] Technically speaking, they show that UEOPL contains the problem of computing an exact fixed point of contraction maps specified by LinearFIXP arithmetic circuits, where the unique fixed point is guaranteed to be rational. For our more general version $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$, the unique fixed point of the underlying map may be irrational., a subclass of CLS to capture problems with a unique solution. It is not known that $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$ is complete for any ${\mathsf{TFNP}}$ class. Notably, to the best of our knowledge, for the known fixed point problems that are complete for some ${\mathsf{TFNP}}$ class, their query complexity in the black-box model is exponential. Examples of such well-known problems include ${\mathsf{PPAD}}$-complete problems Brouwer and Sperner [Pap94, CD09], ${\mathsf{PPA}}$-complete problems Borsuk–Ulam, Tucker [Pap94, ABB20] and MöbiusSperner [DEF+21], CLS-complete problems KKT [FGHS23] and MetricBanach [^6] MetricBanach refers to the problem of computing an approximate fixed point of a contraction map where the distance function $d$ is also part of the input. [DTZ18], and UEOPL-complete problem OPDC [FGMS20].
@@
However, our results indicate that $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$ is dramatically different from all these fixed point problems above in terms of query complexity. Thus, we would like to interpret our results as evidence supporting that $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$ under white-box model might be computationally tractable. Ideally, if it is in ${\mathsf{FP}}$, it would imply many breakthroughs in the fields of verification, semantics, learning theory, and game theory as we discussed before.
@@ 1.1 Sketch of the Main Algorithm
We give a high-level sketch of the main query algorithm for Theorem 1. We start by discretizing the search space. Let $g:[0,n]^{k}\mapsto[0,n]^{k}$ with $g(x):=n\cdot f(x/n)$ and $n:=\lceil 16/(\gamma\varepsilon)\rceil$. It is easy to show that $g$ remains a $(1-\gamma)$-contraction over $[0,n]^{k}$ and it suffices to find a $(16/\gamma)$-fixed point of $g$. Moreover, by rounding the unique fixed point $x^{*}$ of $g$ to an integer point, we know trivially that at least one integer point $x$ in the grid $[0:n]^{k}$, where $[0:n]:=\{0,1,\ldots,n\}$, satisfies $\|x-x^{*}\|_{\infty}\leq 1$ and it is easy to show that any such point $x$ must be a $(16/\gamma)$-fixed point. So our goal is to find a point $x\in[0:n]^{k}$ that satisfies $\|x-x^{*}\|_{\infty}\leq 1$ query-efficiently.
To this end, we use $\texttt{Cand}^{t}$ to denote the set of $[0:n]^{k}$ that remains possible to be close to the unknown exact fixed point $x^{*}$ of $g$. Starting with $\texttt{Cand}^{0}$ set to be the full grid $[0:n]^{k}$, the success of the algorithm relies on whether we can cut down the size of $\texttt{Cand}^{t}$ efficiently. For this purpose we prove a number of geometric lemmas in Section 3 to give a characterization of the exact fixed point $x^{*}$, which lead to the following primitive used by the algorithm repeatedly:
Given $x\in[0,n]^{k}$, $i\in[k]$ and $\phi\in\{\pm 1\}$, we write $\mathcal{P}_{i}(x,\phi)$ to denote the set of points $y\in[0,n]^{k}$ such that $\phi\cdot(y_{i}-x_{i})=\|y-x\|_{\infty}$, where $\mathcal{P}$ is a shorthand for pyramid. Then after querying a point $a\in[0,n]^{k}$, either $a$ was found to be a $(16/\gamma)$-fixed point (in which case the algorithm is trivially done), or one can find $\phi_{i}\in\{\pm 1\}$ for each $i\in[k]$ such that no point in $\mathcal{P}_{i}(a,\phi_{i})$ can be close (within $\ell_{\infty}$-distance $1$) to $x^{*}$ (in which case we can update $\texttt{Cand}^{t}$ by removing all points in $\cup_{i\in[k]}\mathcal{P}_{i}(a,\phi_{i})$).
Given this, it suffices to show that for any set of points $T\subseteq[0:n]^{k}$ (as $\texttt{Cand}^{t}$), there exists a point $a$ to be queried such that for any $\phi_{i}\in\{\pm 1\}$:
@@
is large relative to $|T|$, which is equivalent (up to a factor of $k$) to showing that there exists a point $a$ and $i\in[k]$ such that
$\min\left(\big{|}T\cap\mathcal{P}_{i}(a,+1)\big{|},\big{|}T\cap\mathcal{P}_{i}(a,-1)\right|\big{)}$ (1)
is large relative to $|T|$. This unfortunately turns out to be not true (see Example 1). However, it turns out that such a point (which we will refer to as a balanced point) always exists if we replace the integer grid $[0:n]^{k}$ by the grid of odd-even points: $\texttt{OE}(n,k)$, where $y\in\texttt{OE}(n,k)$ iff $y\in[0:n]^{k}$ and its coordinates are either all odd or all even. To prove the existence of a balanced point, we construct an infinite sequence of continuous maps $\{f^{t}\}$ that can be viewed as relaxed versions of the search for a balanced point. Using Brouwer’s fixed point theorem, every map $f^{t}$ has a fixed point $p^{t}$ and thus, by the Bolzano–Weierstrass theorem, there must be an infinite subsequence of $\{p^{t}\}$ that converges. Letting $p^{*}$ be the point it converges to, we further round $p^{*}$ to $q^{*}\in\texttt{OE}(n,k)$ and show that the latter is a balanced point in the grid. While we show such a point always exists, the brute-force search to find $q^{*}\in\texttt{OE}(n,k)$ is the reason why our algorithm is not time-efficient.

### v2

@@ Abstract
We give an algorithm for finding an $\varepsilon$-fixed point of a contraction map $f:[0,1]^{k}\mapsto[0,1]^{k}$ under the $\ell_{\infty}$-norm with query complexity $O(k\log(1/\varepsilon))$.
@@ 1 Introduction
A particularly important metric space to study the problem of finding a Banach’s fixed point is the $k$-cube $0,1]^{k}$ with respect to the $\ell_{\infty}$-norm, since many important problems can be reduced to that of finding an $\varepsilon$-fixed point (i.e., $x\in[0,1]^{k}$ satisfying $\|f(x)-x\|_{\infty}\leq\varepsilon$) in a $(1-\gamma)$}-contraction map under the $\ell_{\infty}$-norm. Such problems arise from a variety of fields including stochastic analysis, optimization, verification, semantics, and game theory. For example, the classical dynamic programming models mentioned above (Markov decision processes etc.) involve contraction maps under the $\ell_{\infty}$-norm. Furthermore, the same holds for several well-known open problems that have been studied extensively and are currently not known to be in $\mathsf{P}$}. For instance, Condon’s simple stochastic games (SSGs) [Con92] can be reduced to the problem of finding an $\varepsilon$-fixed point in a $(1-\gamma)$-contraction map over $0,1]^{k}$ under the $\ell_{\infty}$-norm. A similar reduction from [EY10] extends to an even broader class of games, namely, Shapley’s stochastic games [Sha53], which lay the foundation of multi-agent reinforcement learning [Lit94]. The same holds also of course for other problems known to be subsumed by SSGs, like parity games, which are important in verification (see e.g. [EJ91, CJK+22]), and mean payoff games [ZP96]. Crucially, in all these reductions, both the approximation parameter $\varepsilon$ and the contraction parameter $\gamma$ are inversely exponential in the input size. Therefore, efficient algorithms in this context are those with a complexity upper bound that is polynomial in $k,\log(1/\varepsilon)$ and $\log(1/\gamma)$.
In this paper we consider general algorithms that access the contraction map in a black-box manner (as an oracle), and study the query complexity of finding an $\varepsilon$-fixed point of a $(1-\gamma)$-contraction map over the $k$-cube $0,1]^{k}$ under the $\ell_{\infty}$-norm (which we denote by $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$). An algorithm under this model is given $k$, $\varepsilon$, $\gamma$, and oracle access to an unknown $(1-\gamma)$-contraction map $f$ over $0,1]^{k}$. In each round the algorithm can send a point $x\in[0,1]^{k}$ to the oracle to reveal its value $f(x)$. The goal of the algorithm is to find an $\varepsilon$-fixed point with as few queries as possible.
@@
There is an $O(k\log({1}/{\varepsilon}))$-query algorithm for $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$.
@@
$\varepsilon/2\geq\|g(x)-x\|_{\infty}=\|(1-\varepsilon/2)f(x)-x\|_{\infty}\geq\|f(x)-x\|_{\infty}-\varepsilon/2.$
@@
In Section 4, we give an $O(k\log(1/\varepsilon))$-query algorithm for $\textsc{Contraction}_{\infty}(\varepsilon/2,\varepsilon/2,k)$, from which Theorem 1 follows. Indeed, note that Observation 1 holds even if $f$ is a non-expansive map (i.e., $f$ has Lipschitz constant $1$: $\|f(x)-f(y)\|_{\infty}\leq\|x-y\|_{\infty}$ for all $x,y\in[0,1]^{k}$). As a result, the same query upper bound applies to $\textsc{NonExp}_{\infty}(\varepsilon,k)$, the problem of finding an $\varepsilon$-fixed point in a non-expansive map over $0,1]^{k}$ under the $\ell_{\infty}$-norm:
@@
There is an $O(k\log({1}/{\varepsilon}))$-query algorithm for $\textsc{NonExp}_{\infty}(\varepsilon,k)$.
@@
There is an $O(k\log({1}/({\varepsilon\gamma})))$-query algorithm for $\textsc{StrContraction}_{\infty}(\varepsilon,\gamma,k)$.
In sharp contrast with Corollary 2, however, we show that it is impossible to strongly approximate an exact fixed point in a non-expansive map over $0,1]^{2}$ under the $\ell_{\infty}$ norm.
@@
The impossibility of strong approximation for expansive maps under the $\ell_{\infty}$ norm, i.e. maps with Lipschitz constant $L>1$, was shown previously by [SS03], and it was conjectured that the same fact should hold also for non-expansive maps ($L=1$). For the $\ell_{2}$ norm, impossibility of strong approximation for non-expansive maps was shown earlier [Sik01].
@@
For any promise problem, one can define a corresponding total search problem, where the black-box can be any function $f$ on the domain and the problem is to compute either a solution or a violation of the promise. In our case, the corresponding total search problem, denoted $\textsc{T-Contraction}_{\infty}(\varepsilon,\gamma,k)$, is the problem of computing for a given function $f:[0,1]^{k}\mapsto[0,1]^{k}$ either an $\varepsilon$-fixed point or a violation of the contraction property, i.e. a pair of points $x,y\in[0,1]^{k}$ such that $\|f(x)-f(y)\|_{\infty}>(1-\gamma)\|x-y\|_{\infty}$. For any promise problem, the corresponding total search problem is clearly at least as hard as the promise problem. For some problems it can be strictly harder (and it may depend on the type of violation that is desired). However, we show that in our case the two versions have the same query complexity.
@@
theoremtotalversion There is an $O(k\log({1}/{\varepsilon}))$-query algorithm for $\textsc{T-Contraction}_{\infty}(\varepsilon,\gamma,k)$.
@@
Remark. It is also important to note that while our algorithm in Theorem 1 is query-efficient, we do not currently see a way of making it time-efficient. The algorithm guarantees that within polynomial queries we can find an $\varepsilon$-fixed point, but each iteration requires a brute force procedure to determine the next query point. We will explain more details of techniques in Section 1.1.
@@
Other Related Work. We have already mentioned the most relevant works addressing the query complexity of computing the fixed point of a contraction map. For continuous functions $f:[0,1]^{k}\mapsto[0,1]^{k}$ that have Lipschitz constant greater than 1 (i.e. are expansive), there are exponential lower bounds on the query complexity of computing a (weak) approximate fixed point [HPV89, CD08]. In a succinct black-box model, it is known that the query complexity of $\mathsf{TFNP}$} problems is polynomial in the description size of the object in the box [KNY19]. Their results do not apply to our setting. For instance, in the $\textsc{T-Contraction}_{\infty}(\varepsilon,\gamma,k)$ problem, the objects are essentially all continuous functions over $0,1]^{k}$, whose description size far exceeds an exponential function of $k$.
@@
$\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$ when considered in the white-box model [^4] The white-box model refers to the model where the function is explicitly given by a polynomial-size circuit, in contrast to the black-box model where the function can only be accessed via an oracle as we studied in this paper. When we talk about a computational problem under the white-box model, we measure the efficiency by the time complexity. can be formulated as a total search problem so that it lies in the class $\mathsf{TFNP}$}. In fact, it is one of the motivating problems in [DP11] to define the class CLS for capturing problems that lie in both $\mathsf{PLS}$} [JPY88] and $\mathsf{PPAD}$} [Pap94]. Later, the problem of computing the exact fixed point of Piecewise-Linear contraction maps, an easier variant of $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$, was placed in UEOPL [FGMS20], a subclass of CLS to capture problems with a unique solution. It is not known that $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$ is complete for any $\mathsf{TFNP}$} class. Notably, to the best of our knowledge, for the known fixed point problems that are complete for some $\mathsf{TFNP}$} class, their query complexity in the black-box model is exponential. Examples of such well-known problems include $\mathsf{PPAD}$}-complete problems Brouwer and Sperner [Pap94, CD09], $\mathsf{PPA}$}-complete problems Borsuk–Ulam, Tucker [Pap94, ABB20] and MöbiusSperner [DEF+21], CLS-complete problems KKT [FGHS23] and MetricBanach [^5] MetricBanach refers to the problem of computing an approximate fixed point of a contraction map where the distance function $d$ is also part of the input. [DTZ18], and UEOPL-complete problem OPDC [FGMS20].
@@
However, our results indicate that $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$ is dramatically different from all these fixed point problems above in terms of query complexity. Indeed, it is a necessary step towards obtaining a polynomial-time general-purpose (i.e., black-box) algorithm for $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$, which, ideally if true, would imply many breakthroughs in the fields of verification, semantics, learning theory, and game theory as we discussed before.
@@ 1.1 Sketch of the Main Algorithm
We give a high-level sketch of the main query algorithm for Theorem 1. We start by discretizing the search space. Let $g:[0,n]^{k}\mapsto[0,n]^{k}$ with $g(x):=n\cdot f(x/n)$ and $n:=\lceil 16/(\gamma\varepsilon)\rceil$. It is easy to show that $g$ remains a $(1-\gamma)$-contraction over $0,n]^{k}$ and it suffices to find a $(16/\gamma)$-fixed point of $g$. Moreover, by rounding the unique fixed point $x^{*}$ of $g$ to an integer point, we know trivially that there exists at least one point $x$ in $\texttt{EVEN}(n,k)$ [^6] Looking ahead, the usage of $\texttt{EVEN}(n,k)$ will guarantee that all queried points lie in the normal grid $\left\{0,\cdots,n\right\}^{k}$, which implies that the approximate fixed point that we will find lies in the normal grid $\left\{0,\cdots,n\right\}^{k}$ as well., which is the set of integral points in the grid $\{0,1,\ldots,n\}^{k}$ with all coordinates being even, that satisfies $\|x-x^{*}\|_{\infty}\leq 1$. Furthermore, it is easy to show that any such point $x$ must be a $(16/\gamma)$-fixed point. So our goal is to find a point $x\in\texttt{EVEN}(n,k)$ that satisfies $\|x-x^{*}\|_{\infty}\leq 1$ query-efficiently.
To this end, we use $\texttt{Cand}^{t}$ to denote the set of $\texttt{EVEN}(n,k)$ that remains possible to be close to the unknown exact fixed point $x^{*}$ of $g$. Starting with $\texttt{Cand}^{0}$ set to be the full set $\texttt{EVEN}(n,k)$, the success of the algorithm relies on whether we can cut down the size of $\texttt{Cand}^{t}$ efficiently. For this purpose we prove a number of geometric lemmas in Section 3 to give a characterization of the exact fixed point $x^{*}$, which lead to the following primitive used by the algorithm repeatedly:
Given $x\in[0,n]^{k}$, $i\in[k]$ and $\phi\in\{\pm 1\}$, we write $\mathcal{P}_{i}(x,\phi)$ to denote the set of points $y\in[0,n]^{k}$ such that $\phi\cdot(y_{i}-x_{i})=\|y-x\|_{\infty}$, where $\mathcal{P}$ is a shorthand for pyramid; see Figure 1 for illustrations. Then after querying a point $a\in[0,n]^{k}$, either $a$ was found to be a $(16/\gamma)$-fixed point (in which case the algorithm is trivially done), or one can find $\phi_{i}\in\{\pm 1\}$ for each $i\in[k]$ such that no point in $\mathcal{P}_{i}(a,\phi_{i})$ can be close (within $\ell_{\infty}$-distance $1$) to $x^{*}$ (in which case we can update $\texttt{Cand}^{t}$ by removing all points in $\cup_{i\in[k]}\mathcal{P}_{i}(a,\phi_{i})$).
Given this, it suffices to show that for any set of points $T\subseteq\texttt{EVEN}(n,k)$ (as $\texttt{Cand}^{t}$), there exists a point $a$ to be queried such that for any $\phi_{i}\in\{\pm 1\}$:
@@
is large relative to $|T|$. Quantitatively, we show that there exists a point $a$ such that for any $\phi_{i}\in\{\pm 1\}$, the set $T\cap\left(\bigcup_{i\in[k]}\mathcal{P}_{i}(a,\phi_{i})\right)$ has size at least half of $|T|$. This bound is clearly sharp and we refer to such point $a$ as balanced point.
To prove the existence of a balanced point, we construct an infinite sequence of continuous maps $\{f^{t}\}$ that can be viewed as relaxed versions of the search for a balanced point. Using Brouwer’s fixed point theorem, every map $f^{t}$ has a fixed point $p^{t}$ and thus, by the Bolzano–Weierstrass theorem, there must be an infinite subsequence of $\{p^{t}\}$ that converges. Letting $p^{*}\in[0,n]^{k}$ be the point it converges to, we can show that $p^{*}$ must be a desired balanced point. We could further round $p^{*}$ to $q^{*}\in\left\{0,\dots,n\right\}^{k}$ and show that the latter is a balanced point in the grid. While we show such a point always exists, the brute-force search to find $q^{*}\in\left\{0,\cdots,n\right\}^{k}$ is the reason why our algorithm is not time-efficient.
$x$
(a) The light blue region is given by $\mathcal{P}_{1}(x,-1)$ where $x$ is the center point.
$x$
(b) The light blue region is given by $\mathcal{P}_{1}(x,-1)$ where $x$ is the center point.
Figure 1: Pyramids in two-dimensional and three-dimensional cubes.